a fluid with a vapor 0.4 pascal and the specific gravity of 14 is used in a barometer. the fluids column height is 0.8m.find the atmospheric pressure in kpa


1
Expert's answer
2021-11-16T19:21:48-0500

Density of the fluid "\rho" = specific gravity "\times" density of water "=14\times 1000=14000kg/m^3"

Let "P_v=" vapor pressure=0.4.

From the principle of barometer, we have "\rho gh=P_{atm}-P_v"

"P_{atm}=\rho gh +P_v=14000\times 9.8\times 0.8 +0.4=109760.4Pa"

"P_{atm}=109.7604kpa"